OpenStudy (anonymous):

i am got 57 in term 1 that is worth 20 % . i got 61 in term 2 that is worth 40 %. term 3 is worth 40 %. what is the minimum i need to pass this course?

OpenStudy (anonymous):

57 * .2 = 11.4 61 * .4 = 24.4 11.4+24.4 = 35.8 you need a 60 to pass, so 60-35.8 = 24.2 you need x * .4 = 24.2 to pass term 3 x * .4 = 24.2 divide both sides by .4 60.5 you need at least a 60.5% in term 3 to pass.

OpenStudy (anonymous):

i forgot to add a pass is 50

OpenStudy (anonymous):

ahh okay so 50-35.8 = 14.2 x * .4 = 14.2 x=35.5 you need at least a 35.5% to pass.

OpenStudy (anonymous):

thank you sooo much i feel alot more better ty